Include formula for excel
f1-5 (L01,2) (Depreciation Computations-Four Methods) Robert Parish Corporation purchased a new machine for its E11 assembl a salvage value of $12,900 at the end of its service life. Its life is estimated at 5 years,and its working hours are estimated at y process on August 1, 2017. The cost of this machine was $117,900. The company estimated that the machine would ave 21,000 hours. Year-end is December 31. Instructions Compute the depreciation expense under the following methods. Each of the following should be considered unrelated. (a) Straight-line depreciation for 2017. (b) Activity method for 2017, assuming that machine usage was 800 hours. d Sum-of-the-years-digits for 2018. d) Double-declining-balance for 2018.
Answer #1

a. Depreciation 2017 for 5 months

Depreciation amount = (original cost - salvage value)

= 105,000

= [105,000*5]5/12

= 218,750

b. Depreciation for 2017

= [105,000/21,000]*800

=4,000

c. Sum of year digit = (5+4+3+2+1)=15

Sum of year digit depreciation for 2018

= (105,000*5/15*7/12)+(105,000*4/15*5/12)

= 32,083

d. Depreciation under double declining balance method = (cost - accumulated depreciation) / estimated useful life * 2

Straight line depreciation = 105,000/5

= 21,000

Depreciation % = 21,000/105,000

= 20%

Double declining % = 20%*2 = 40%

Depreciation for 2017 = (117,900*40%)*5/12

= 19,650

Book value at beginning of year 2

= 117,900 - 19,650

= 98,250

Depreciation for 2018 = 98,500*40%

= 39,300